# Low-Rank Tensor Decomposition over Finite Fields

### v2

@@ 2.1 Runtime
In our algorithm, we first factor $\overline{T}=M\mathrm{rref}\left(\overline{T}\right)$ and construct $T^{\prime}$ via Gaussian elimination. Since $J=\mathrm{rank}\left(\overline{T}\right)$ must be at most $R$ in order for $T$ to have a rank-$R$ decomposition, we can terminate early if we detect that this condition fails, yielding a running time of $O(Rn^{3})$444 Given a $n\times m$ matrix $A$, we can construct $C,F$ s.t. $A=CF$, $F=\mathrm{rref}\left(A\right)$, and $C$ is invertible, as follows: 1. initialize $(C,F)=(I_{n},A)$ and $r=0$: 2. for $j$ from 00 to $m-1$ inclusive: (a) find some $i\geq r$ s.t. $F_{i,j}\neq 0$ (b) if such $i$ does not exist: skip to the next iteration of the for-loop (c) else: i. swap rows $i$ and $r$ of $F$, and swap columns $i$ and $r$ of $C$ ii. for $k=r+1,\dots n-1$: A. let $s=F_{k,j}$ B. set $F_{k,:}$ to $F_{k,:}-sF_{r,:}$ C. set $C_{:,r}$ to $C_{:,r}+sC_{:,k}$ (here $C_{:,r}$ denotes column $r$ of $C$) iii. increment $r$ The rank of $A$ is exactly the number of times we reach (c). If we only need to return $(C,F)$ under the condition $\mathrm{rank}\left(A\right)\leq R$, for some rank limit $R$, and return a failure statement otherwise, we can terminate early if we ever detect that (c) has been reached more than $R$ times. Thus, this procedure performs $O(Rn(m+n))$ operations. In our main algorithm, $A=\overline{T}$ has shape $n\times n^{2}$, so this procedure runs in $O(Rn^{3})$ time. .
Then we enumerate every possible $U$, of which there are $\leq|\mathbb{F}|^{JR}$ of them. For each $U$, we perform Gaussian elimination on the matrix $\Xi$, which has dimensions $J\times(R+1)$ and has each element in its rightmost column being a $n\times n$ matrix, so we incur $O\left(R^{3}n^{2}\right)$ runtime.
We then ensure there is no $r$ s.t. $\Omega_{r,:}=\overrightarrow{0}$ and $D_{r}\neq O$, which takes $O(J(R+n^{2}))$ time. If this check passes, we proceed to the rref case analysis, during which we compute the ranks of one or more matrices $D_{i}$, each of which takes $O(n^{3})$ time via Gaussian elimination.
If $R=3$ we also have to enumerate all rank-2 decompositions of a rank-2 $n\times n$ matrix $D_{i}$ in the worst case, which takes $O(n^{3}+|\mathbb{F}|^{4}n)$ time (if we precompute all invertible $2\times 2$ matrices over $\mathbb{F}$).
If we pass the case analysis, we construct $B$ and $C$ in $O(n^{3})$ time, then $A:=\left[\sum_{0\leq j<J}M_{i,j}U_{r,j}\right]_{r,i}$ in $O(RJn)$ time, and terminate.
Since we are guaranteed that $J\leq R$ (as otherwise no solution exists), our final runtime is
$O\left(Rn^{3}+\underbrace{|\mathbb{F}|^{R^{2}}}_{\textrm{fix }U}\left(\underbrace{R^{3}n^{2}}_{\textrm{row-reduce }\Xi}+\underbrace{R(n^{2}+R)}_{\textrm{check rows of }\Xi}+Rn^{3}+\left(\begin{cases}n^{3}+|\mathbb{F}|^{4}n&R=3\\ 0&R\leq 2\end{cases}\right)\right)+n^{3}+R^{2}n\right)\\ =O\left(|\mathbb{F}|^{R^{2}}\left(R^{3}n^{2}+Rn^{3}+\left(\begin{cases}n^{3}+|\mathbb{F}|^{4}n&R=3\\ 0&R\leq 2\end{cases}\right)\right)\right)$.
Thus, for any fixed $\mathbb{F}$ and fixed $R$ our algorithm runs in $O(n^{3})$ time, with hidden constant $f(|\mathbb{F}|,R)$ polynomial in $|\mathbb{F}|$ and exponential in $R$; for a loose bound we can set $f(|\mathbb{F}|,R)=|\mathbb{F}|^{R^{2}+4}R^{3}$. For $\mathbb{F}=\mathbb{Z}/2\mathbb{Z}$ and $R=3$, the coefficient on the $n^{3}$ term is $\approx 2^{3^{2}}\cdot(3+1)=4\cdot 2^{9}=2048$.
2.2 Beyond rank 3
We do not know how to make our algorithm run in $\mathrm{poly}(n)$ time for $R=4$. Consider the case $E=\begin{bmatrix}1&&w&x\\ &1&y&z\end{bmatrix}$, $\mathrm{rank}\left(D_{0}\right)=2,\ \mathrm{rank}\left(D_{1}\right)=2$: unlike when $E=\begin{bmatrix}1&&x\\ &1&y\end{bmatrix}$, we cannot simply force $B_{3,:},C_{3,:}=\overrightarrow{0}$, as the system of equations might become too restrictive. One might also notice that $D_{0}+D_{1}=B_{0,:}\times C_{0,:}+B_{1,:}\times C_{1,:}$, so $D_{0}+D_{1}$ has rank $\leq 2$; however, if it happens to have rank 1, then there still might be exponentially many solutions for $(B_{0,:},C_{0,:},B_{1,:},C_{1,:})$, so it is too expensive to simply iterate over all such solutions and check for feasibility of $(B_{2,:},C_{2,:},B_{3,:},C_{3,:})$.
@@ 3 With wildcards
Over arbitrary finite fields, there is an issue of scale invariance, i.e., each outer product $a\times b\times c$ can be replaced with $(\alpha a)\times(\beta b)\times(\gamma c)$ for any $\alpha,\beta,\gamma$ s.t. $\alpha\beta\gamma=1$, allowing any one tensor decomposition to be transformed into many other tensor decompositions that evaluate to the same tensor. This scaling freedom makes constructing gadgets for NP-hardness less obvious.
@@ 3.1 Rank 1 with wildcards
Here we outline a polynomial-time algorithm for rank-1 decomposition with wildcards over a fixed finite field $\mathbb{F}$.
First, we rewrite the problem as follows: given an arbitrary subset $S\subseteq\left\{0,\dots n-1\right\}^{3}$ and set of elements $T_{i,j,k}\in\mathbb{F}\ \forall(i,j,k)\in S$, find variables $a_{0\dots n-1},b_{0\dots n-1},c_{0\dots n-1}$ s.t. $a_{i}b_{j}c_{k}=T_{i,j,k}\ \forall(i,j,k)\in S$.
@@
Otherwise, we proceed with the observation that all the equations $a_{i}b_{j}c_{k}=T_{i,j,k}$ involve pure products, i.e., expressions of the form $\left(\prod_{i}a_{i}^{p_{i}}\right)\left(\prod_{j}b_{j}^{q_{j}}\right)\left(\prod_{k}c_{k}^{r_{k}}\right)$, for integers $p_{i},q_{j},r_{k}$. If we allow arbitrary multiplication and division between such expressions, all resulting expressions will still be of this form, although some $p_{i},q_{j},r_{k}$ might be negative.
This observation suggests that we can solve this problem somewhat like a linear system of equations: construct a $|S^{\prime}|\times 3n$ integer matrix $E$ and a $|S^{\prime}|\times 1$ $\mathbb{F}$-vector $W$ s.t. each equation $a_{i}b_{j}c_{k}=T_{i,j,k}$ for $(i,j,k)\in S^{\prime}$ is encoded as $\prod_{m}v_{m}^{E_{l,m}}=W_{l}$ for some $l$, where the variables have been re-indexed as $v_{0,\dots n-1}=a_{0,\dots n-1},\ v_{n,\dots 2n-1}=b_{0,\dots n-1},\ v_{2n,\dots 3n-1}=c_{0,\dots n-1}$. Since we restrict ourselves to $S^{\prime}$, each $W_{l}$ is nonzero. During row reductions, elements in $W$ have to be treated differently: when adding between two rows, multiply their values in $W$; when scaling a row by $s$, raise its value in $W$ to the $s$-th power.
However, once we reduce $\left[\begin{array}[]{c|c}E&W\end{array}\right]$ to some normal form, elements in $E$ might not be ±$\pm 1$, so extracting a solution for the variables $v_{i}$ may involve computing $r$-th roots in $\mathbb{F}$. An explicit example of this phenomenon is the tensor
$T=\begin{bmatrix}\begin{bmatrix}1&?\\ ?&1\end{bmatrix},\begin{bmatrix}?&2\\ 1&?\end{bmatrix}\end{bmatrix},$
where ? represents wildcards (equivalently, $S=\left\{(0,0,0),(1,1,0),(0,1,1),(1,0,1)\right\}$, $T_{0,0,0}=1,\ T_{1,1,0}=1,\ T_{0,1,1}=2,\ T_{1,0,1}=1$). Solving this instance is equivalent to solving the equations
$\left\{a_{0}b_{0}c_{0}=1,\ a_{1}b_{1}c_{0}=1,\ a_{0}b_{1}c_{1}=2,\ a_{1}b_{0}c_{1}=1\right\};$
by combining these equations, we get $\frac{(a_{0}b_{1}c_{1})(a_{1}b_{0}c_{1})}{(a_{0}b_{0}c_{0})(a_{1}b_{1}c_{0})}=\frac{2\cdot 1}{1\cdot 1}\Rightarrow\left(\frac{c_{1}}{c_{0}}\right)^{2}=2$, so solving this equation requires us to compute a square root of 2.
To resolve this issue, we use the fact that the multiplicative group of any finite field is cyclic, so if $\mathbb{F}$ is fixed we can find a primitive element (i.e., multiplicative generator) $\mu$ via brute force and define $u_{i},w_{i}$ s.t. $v_{i}=\mu^{u_{i}},\ W_{i}=\mu^{w_{i}}$; then our problem reduces to solving the linear system of equations $Eu=w$ for the vector $u$ over $\mathbb{Z}/(|\mathbb{F}|-1)\mathbb{Z}$.
@@
From [6], solving the matrix-vector equation $xA=b$ over $\mathbb{Z}/N\mathbb{Z}$ for $n\times m$ $A$ can be done in $O(nm^{\omega-1})$ many operations, where $\omega\leq 2.371552$ [8] is the exponent for matrix multiplication, with all integers throughout the algorithm having magnitude $\leq N$. By transposing our system to $u^{\intercal}E^{\intercal}=w^{\intercal}$ and noting that $E$ has dimensions $|S^{\prime}|\times 3n$, we see that solving our system takes $O(n|S^{\prime}|^{\omega-1})$ operations, which is within $O(n^{3\omega-2})$ since $|S^{\prime}|\leq n^{3}$. For fixed $\mathbb{F}$, since all integers used in this algorithm have magnitude $\leq|\mathbb{F}|-1$, each operation takes $O(1)$ time, so our algorithm runs in polynomial time. If we account for the number of wildcards in $T$ in the time complexity, then by letting $k$ be the number of wildcards, $|S^{\prime}|\leq|S|=n^{3}-k$, so we get $O(n(n^{3}-k)^{\omega-1})$.
@@
The main bottleneck in this algorithm is finding a primitive element of $\mathbb{F}$, for which an efficient algorithm is not known [7]. On the other hand, for $\mathbb{F}=\mathbb{F}_{p^{k}}$, the total input size of a problem instance would be linear w.r.t. $\log(p)$ and $k$ [^6] Every element in $\mathbb{F}_{p^{k}}$ is a linear combination of $1,\alpha,\dots,\alpha^{k-1}$ with coefficients in $\mathbb{Z}/p\mathbb{Z}$, for some element $\alpha$, so every element can be specified with a $k$-element vector over $\mathbb{Z}/p\mathbb{Z}$.. Because of this, we are not sure what the computational hardness of rank-1 wildcard decomposition is w.r.t. input size when $p,k$ are not fixed parameters.

### v3

@@ 2.3 Runtime
To provide a good upper bound on the asymptotic time complexity of our algorithm, we take advantage of the fact that all matrices we deal with must have low rank in order for a solution to the original tensor decomposition problem to exist. For this, we utilize an efficient algorithm that returns full-rank factorizations for matrices of rank below some threshold and returns nothing otherwise.
Lemma 5.
Given a matrix $M\in\mathbb{F}^{m\times n}$ and integer $r^{\prime}$, we can do the following in $O(mn\mathrm{min}\left(r^{\prime},\mathrm{rank}\left(M\right)\right))$ time:
  • If $\mathrm{rank}\left(M\right)\leq r^{\prime}$: return some $C\in\mathbb{F}^{m\times\mathrm{rank}\left(M\right)}$ and $F\in\mathbb{F}^{\mathrm{rank}\left(M\right)\times n}$ s.t. $M=CF$
  • Else: return nothing
Proof.
We run Gaussian elimination with early termination:
  1. 1.
    initialize $(C,F)\leftarrow(I_{m},M)$ and $r\leftarrow 0$:
  2. 2.
    for $j$ from 00 to $n-1$, inclusive:
    1. (a)
      if there is no $r\leq i<m$ s.t. $F_{i,j}\neq 0$: continue to the next iteration of the loop
    2. (b)
      else:
      1. i.
        find some $r\leq i<m$ s.t. $F_{i,j}\neq 0$
      2. ii.
        swap rows $i$ and $r$ of $F$
      3. iii.
        swap columns $i$ and $r$ of $C$
      4. iv.
        let $\sigma=F_{r,j}$
      5. v.
        set $F_{r,:}\leftarrow F_{r,:}/\sigma$
      6. vi.
        set $C_{:,r}\leftarrow\sigma C_{:,r}$
      7. vii.
        for $k$ from $r+1$ to $m-1$, inclusive:
        1. A.
          let $s=F_{k,j}$
        2. B.
          set $F_{k,:}\leftarrow F_{k,:}-sF_{r,:}$
        3. C.
          set $C_{:,r}\leftarrow C_{:,r}+sC_{:,k}$
      8. viii.
        increment $r$
      9. ix.
        if $r>r^{\prime}$: terminate and return nothing
  3. 3.
    return $(C_{:,:r},F_{:r,:})$
If there was no early termination (ix.), this algorithm would transform $F$ into row echelon form (not necessarily reduced row echelon form), so it would reach step (b) exactly $\mathrm{rank}\left(M\right)$ many times. Thus, with early termination, the algorithm will reach step (b) $\mathrm{min}\left(r^{\prime},\mathrm{rank}\left(M\right)\right)$ many times, correctly terminating if $\mathrm{rank}\left(M\right)>r^{\prime}$.
Finally, each execution of step (b) involves $O(m)$ many row operations in $F$ and column operations in $C$, each of which takes $O(n)$ time, so the total runtime is $O(mn\mathrm{min}\left(r^{\prime},\mathrm{rank}\left(M\right)\right))$. 777 As a sidenote, we can slightly modify this algorithm to solve Lemma 1 in $O(mn\mathrm{rank}\left(M\right))$ time: remove the early termination (ix.); maintain another matrix $T$ that is initialized to $I_{m}$, and perform the same operations on it as the ones we do to $F$ throughout the algorithm;
In our algorithm, we first extract a basis of tensor slices $T_{i,:,:}$ by row-reducing the $n\times n^{2}$ matrix $\widetilde{T}:=\begin{bmatrix}\vdots\\ \hline\cr\overrightarrow{T_{i,:,:}}\\ \hline\cr\vdots\end{bmatrix}_{i}$. We also check whether $\mathrm{rank}\left(\widetilde{T}\right)\leq R$ and immediately exit if this condition is not satisfied. Using Lemma 5, we can complete this entire step in $O(Rn^{3})$ time.
Then, for each possible $A^{\prime}$, we solve a system of linear equations over $R$ many matrices $M_{r}$ of rank $\leq 1$. Since $A^{\prime}$ has $R$ columns and $\leq R$ rows, there are $O(|\mathbb{F}|^{R^{2}})$ many of them.
To solve the system over $M_{r}$, we first separate it into independent parts, of which there are $\leq R$ many. Then we solve each part by enumerating full-rank factorizations of a matrix with rank $\leq g(R)$, for some function $g$, and compute $O(R)$ many additional rank factorizations to extract $B_{r,:}$ and $C_{r,:}$ if a solution is found.
To figure out $g$:
  • If $R\leq 2$: all coefficient matrices $E$ are separable into single rows, so we do not enumerate full-rank factorizations. For this, we define $g(R):=0$.
  • If $R=3$: the most difficult case occurs when $E=\begin{bmatrix}1&&x\\ &1&y\end{bmatrix}$ and at least one of $D_{0}$ or $D_{1}$ has rank 2.
  • If $R=4$: the most difficult case occurs when $E=\begin{bmatrix}1&&w&x\\ &1&y&z\end{bmatrix}$ and at least one of $\Phi:=\left[\begin{array}[]{c|c}D_{0}&D_{1}\end{array}\right]$ or $\Theta:=\left[\begin{array}[]{c}D_{0}\\ \hline\cr D_{1}\end{array}\right]$ has rank 4.
    • We also have the subcase where $\mathrm{rank}\left(D_{0}\right)=\mathrm{rank}\left(D_{1}\right)=2\ \land\ \mathrm{rank}\left(\Phi\right)=\mathrm{rank}\left(\Theta\right)=3\ \land\ \mathrm{rank}\left(U\right)=3$, where we enumerate every $3\times 4$ matrix $F$ and solve for row-vectors $v_{i}$ in a linear system. Since this system contains $O(n)$ total scalar variables, it can be calculated in $O(n^{3})$ time, so this entire subcase can be solved in $O(|\mathbb{F}|^{12}n^{3})$ time, which is asymptotically less costly than the case where at least one of $\Phi$ or $\Theta$ has rank 4.
Thus, we have $g(1)=0,\ g(2)=0,\ g(3)=2,\ g(4)=4$.
Since all rank factorizations are done for $O(n)\times O(n)$ matrices, computing each one takes $O(n^{3})$ time, so we can upper bound the total time complexity of our algorithm at $O\left(Rn^{3}+|\mathbb{F}|^{R^{2}}R\left(|\mathbb{F}|^{g(R)^{2}}+R\right)n^{3}\right)=O\left(\left(R+|\mathbb{F}|^{R^{2}}R\left(|\mathbb{F}|^{g(R)^{2}}+R\right)\right)n^{3}\right)$, which is sufficient to prove Theorem 1.
2.3.1 Tighter analysis
By analyzing the rank factorizations of our algorithm more carefully, we can obtain a tighter bound.
Consider enumerating full-rank factorizations for a $O(n)\times O(n)$ matrix $M$ with rank $r$. We first compute an initial factorization $M=CF$, where $C$ has $r$ columns and $F$ has $r$ rows, which can be done in $O(rn^{2})$ time. Then, for each $X\in\mathrm{GL}\left(r,\ \mathbb{F}\right)$, we compute $(CX,X^{-1}F)$, which takes $O(nr^{2})$ time (we can precompute each $X^{-1}$ ahead of time). However, we may also have to set each appropriate variable matrix $M_{r}$ to a summand $(CX)_{:,r^{\prime}}(X^{-1}F)_{r^{\prime},:}$ of the rank factorization, e.g., in the rows-sharing-one-column case, in order to continue with the rest of the algorithm, so constructing these $M_{r}$ will take additional $O(rn^{2})$ time. Thus, the total time needed to enumerate all full-rank factorizations and set the appropriate $M_{r}$ is $O(|\mathbb{F}|^{r^{2}}(rn^{2}+r^{2}n))$.
Lastly, we must address the one subcase under $R=4$ that we mentioned earlier: where $\mathrm{rank}\left(D_{0}\right)=\mathrm{rank}\left(D_{1}\right)=2\ \land\ \mathrm{rank}\left(\Phi\right)=\mathrm{rank}\left(\Theta\right)=3\ \land\ \mathrm{rank}\left(U\right)=3$, and we must solve $F\left[\begin{array}[]{c|c}v_{0}&\overrightarrow{0}\\ \hline\cr\overrightarrow{0}&v_{1}\\ \hline\cr v_{2}&v_{2}\\ \hline\cr v_{3}&zv_{3}\end{array}\right]=W_{*}$, where $F\in\mathbb{F}^{3\times 4}$, $v_{i}\in\mathbb{F}^{1\times n}$, and $W_{*}\in\mathbb{F}^{3\times 2n}$ forms the right side of a rank-3 factorization of $\Phi\in\mathbb{F}^{n\times 2n}$ (i.e. $\Phi=U_{*}W_{*}$ for some $U_{*}\in\mathbb{F}^{n\times 3}$). We can actually separate this system into $n$ many disjoint linear systems $F\begin{bmatrix}v_{0,i}&0\\ 0&v_{1,i}\\ v_{2,i}&v_{2,i}\\ v_{3,i}&zv_{3,i}\end{bmatrix}=\left[\begin{array}[]{c|c}W_{:,i}&W_{:,n+i}\end{array}\right]$ for each $0\leq i<n$, where each system only consists of $8$ variables instead of $O(n)$. After accounting for the enumeration of $F$ and the factorization $\Phi=U_{*}W_{*}$, we can actually solve this entire subcase in $O(3n^{2}+|\mathbb{F}|^{12}8^{3}n)$ time; this is less than the other subcase under $R=4$ where $\mathrm{rank}\left(\Phi\right)=4\ \lor\ \mathrm{rank}\left(\Theta\right)=4$, which requires $O(|\mathbb{F}|^{16}(4n^{2}+4^{2}n))$ time.
Applying these findings to our algorithm, we enumerate rank factorizations of a $O(n)\times O(n)$ matrix of rank $g(R)$; and if a solution is found, we rank-factor the $D_{i}$ matrices, each of which must have rank $\leq R$ by the rank precondition (or else no solution exists and we terminate), which takes $O(R)\cdot O(Rn^{2})=O(R^{2}n^{2})$ additional time.
Thus, the time complexity of our algorithm is $O\left(Rn^{3}+|\mathbb{F}|^{R^{2}}R\left(|\mathbb{F}|^{g(R)^{2}}\left(g(R)n^{2}+g(R)^{2}n\right)+R^{2}n^{2}\right)\right)$. Notice that the $n^{3}$ now has a small coefficient compared to the other terms. It may be possible to obtain slightly tighter bounds through more careful analysis.
@@ 3 With wildcards
Over arbitrary finite fields, there is an issue of scale invariance: each outer product $a\times b\times c$ can be replaced with $(\alpha a)\times(\beta b)\times(\gamma c)$ for any $\alpha,\beta,\gamma$ s.t. $\alpha\beta\gamma=1$, which allows a single tensor decomposition to be transformed into many other tensor decompositions that evaluate to the same tensor. This scaling freedom makes constructing gadgets for NP-hardness more difficult.
@@ 3.1 Rank 1 with wildcards
For rank-1 decomposition with wildcards, we give a polynomial-time algorithm over a fixed finite field $\mathbb{F}$.
First, we rewrite the problem as follows: given an arbitrary subset $S\subseteq\left\{0,\dots n-1\right\}^{3}$ of tensor indices and a set of elements $T_{i,j,k}\in\mathbb{F}\ \forall(i,j,k)\in S$, find variables $a_{0\dots n-1},b_{0\dots n-1},c_{0\dots n-1}$ s.t. $a_{i}b_{j}c_{k}=T_{i,j,k}\ \forall(i,j,k)\in S$.
@@
Otherwise, we observe that all the equations $a_{i}b_{j}c_{k}=T_{i,j,k}$ involve multiplications only. This suggests we replace all scalars with exponentials under the same base, as this turns multiplications into additions.
Indeed, it is well-known that the multiplicative group of any finite field is cyclic, so there exists some $\mu\in\mathbb{F}$ s.t. its powers cover all of $\mathbb{F}\setminus\left\{0\right\}$. We can then substitute the variables $a_{i}=\mu^{p_{i}},\ b_{j}=\mu^{q_{j}},\ c_{k}=\mu^{r_{k}}$ and tensor elements $T_{i,j,k}=\mu^{t_{i,j,k}}$ to get the equations $p_{i}+q_{j}+r_{k}\equiv t_{i,j,k}\ (\bmod|\mathbb{F}|-1)\ \forall(i,j,k)\in S^{\prime}$. Thus, solving our problem is equivalent to solving a linear system over $\mathbb{Z}/(|\mathbb{F}|-1)\mathbb{Z}$.
@@
From [6], solving the matrix-vector equation $xA=b$ over $\mathbb{Z}/N\mathbb{Z}$ for $n\times m$ $A$ can be done in $O(nm^{\omega-1})$ many operations, where $\omega\leq 2.371552$ [8] is the exponent for matrix multiplication, with all integers throughout the algorithm having magnitude $\leq N$. Since our system consists of $|S^{\prime}|$ equations over $3n$ variables, it can be solved in $O(n|S^{\prime}|^{\omega-1})$ operations, which is within $O(n^{3\omega-2})$ since $|S^{\prime}|\leq n^{3}$.
@@
The main bottleneck in this algorithm is finding a primitive element of $\mathbb{F}$, for which an efficient algorithm is not known [7]. On the other hand, for $\mathbb{F}=\mathbb{F}_{p^{k}}$, the total input size of a problem instance would be linear w.r.t. $\log(p)$ and $k$  [^9] Every element in $\mathbb{F}_{p^{k}}$ is a linear combination of $1,\alpha,\dots,\alpha^{k-1}$ with coefficients in $\mathbb{Z}/p\mathbb{Z}$, for some element $\alpha$, so every element can be specified with a $k$-element vector over $\mathbb{Z}/p\mathbb{Z}$.. Because of this, we are not sure what the computational hardness of rank-1 wildcard decomposition is when $p,k$ are not fixed.